i know the answer but can't seem to make out an equation! out of my teacher's own problem packet. it reads...
For a Kindergarten party, Mrs. Tully bought four more bags of jellybeans selling at 35 cents a bag than lemon drops selling at 45 cents a bag. She spent $3.00 for the candy. How many bags of jellybeans did she buy???
i know Mr.s Tully bought six bags of jelly beans and two bags of lemon drops but i can't figure out how to write it in a "teacher acceptable" equation.
thanks a ton!!!
~misguided mathmatitian

Answer by mukhopadhyay(490)

Number of bags containing lemon drops = x
=> Number of bags containing jelly beans = x+4
Cost of lemon drops = 45x
Cost of jelly beans = 35(x+4)
Total cost = 300
=> 45x+35(x+4) = 300
=> 45x+35x+140 = 300
=> 80x = 160
=> x=2
Answer: You got the answer correct.